# Yong Ji1, Junye Li1 and Rui Yang∗2,3,4

### v1

@@ 4.1 Variational principle of level sets
denote the level set of $\alpha$ w.r.t. $\varphi$ consisting of the set of points of $X$ such that $\varphi$ is taking constant value $\alpha$ on $V_{f}(x)$. Then continuity of $\varphi$ implies that $M_{f}(\varphi,\alpha)$ is a nonempty closed subset of $M_{f}(X)$.
Condition A: For each given $\varphi$ and $\alpha\in L_{\varphi}$, we always assume that $M_{f}(\varphi,\alpha)$ is a convex subset.
@@
    Let $\Phi:=\{\phi_{n}\}$ be an asymptotically additive potential, i,e. for any $\delta>0$, there exists $\psi_{\delta}\in C(X)$ such that $\limsup_{n\to\infty}\frac{1}{n}||\phi_{n}-S_{n}\psi_{\delta}||<\delta$. We define
@@
    Specially, when sequence $\log\phi_{n}(x)=S_{n}\psi(x)$ is exactly an additive function, where $\psi\in C(X)$ and $S_{n}\psi(x)=\sum_{i=0}^{n-1}\psi(f^{i}(x))$, in this case we have
    $K_{\alpha}=\{x\in X:\forall\mu\in V_{f}(x),\ \int\varphi\mathop{}\!\mathrm{d}\mu=\alpha\},$
    $M_{f}(\varphi,\alpha)=\{\mu\in M_{f}(X):\int\varphi\mathop{}\!\mathrm{d}\mu=\alpha\}.$
Given the nonempty open subset $U\subset X$, we put
@@
Let $(X,d,f)$ be a dynamical system with specification property and $U\subset X$ be a nonempty open subset. Suppose that let $\varphi:M_{f}(X)\to\mathbb{R}$ be a continuous function satisfying Condition A. For any $\alpha\in L$ and $S\in\{B,P\}$, we have
@@
$=$ $\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
@@
Put $G_{\mu}^{\omega,U}:=G_{\mu}\cap U\cap\left\{x\in X:\omega_{f}(x)=X\right\}$. By (3.14), there exists a constant $c>0$ which is not dependent on $\mu$, such that
@@
denote the $\alpha$-level set of $\psi$. If $\alpha\in\mathbb{R}$ is a constant such that $E_{\alpha}\not=\emptyset$, then we have
$\overline{\rm mdim}_{M}^{B}(f,E_{\alpha},d)=\overline{\rm mdim}_{M}^{P}(f,E_{\alpha},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}\xi<\epsilon}h_{\mu}(f,\xi),$
@@
$K_{\alpha}=\{x\in X:\forall\mu\in V_{f}(x),\int\psi\mathop{}\!\mathrm{d}\mu=\alpha\}$
satisfies Condition A. Fix $x\in X$. Since the statement $\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\psi(f^{j}x)=\alpha$ is equivalent to the statement $V_{f}(x)\subset M_{f}(\psi,\alpha)$, this shows $K_{\alpha}=E_{\alpha}$. We get the desired result by Theorem 4.1. ∎
Readers can turn to [O03, Tho09, ZC13] and references therein for the multifractal analysis of level sets using topological pressure. Then by Theorem 4.1 one can get analogous results for the metric mean dimensions of level sets of some special functions determined by the dynamical behaviors of the system.
@@ 4.2 Linking metric mean dimension and mean Li-Yorke chaotic
In this subsection, we link the concepts of metric mean dimension and mean Li-Yorke chaos in the sense of infinite entropy. We give an quantitative characterization for the chaotic phenomenon of infinite entropy systems. Namely, we show the packing metric mean dimension of the set of mean Li-Yorke pairs is bounded from blow by the metric mean dimension of the phase space (Theorem 4.3).
@@
Let $(X,d,f)$ be a dynamical system with specification property with positive topological entropy. Then we have
$\overline{\rm mdim}^{P}_{M}(f\times f,{\rm MLY}(f),d\times d)\geq\overline{\rm mdim}_{M}(f,X,d).$
@@
By the well-known fact that positive entropy implies mean Li-Yorke chaotic (cf.[Dow14]), we can choose two distinct points of $x,y\in X$, a sequence $\{n_{i}\}$ of positive integer that converges to $\infty$ and $\theta>0$ (depending on $x,y$) such that $\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}d(f^{j}x,f^{j}y)>\theta$ for all sufficiently large $i$. We put $\mu_{n_{i}}=\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}\delta_{(f^{j}x,f^{j}y)}$. Then we can choose an accumulation point $\mu_{0}\in M_{f\times f}(X\times X)$ of $\{\mu_{n_{i}}\}$ satisfying the desired claim. ∎
Now take $\nu\in M_{f}(X)$ and set $\mu_{1}\in M_{f\times f}(X\times X)$ defined by $\mu_{1}=\phi_{\ast}\nu$, where $\phi:X\to X\times X$ is the diagonal map given by $\phi(x)=(x,x)$. Then one has $\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{1}=0$. Assume that $\overline{\rm mdim}_{M}(f,X,d)$ is finite. For any $\gamma>0$, by Theorem 2.4 we pick $\epsilon_{k}\to 0$ as $k$ goes to $\infty$ and $\rho_{k}\in M_{f}(X)$ such that for each $k$,
$\frac{1}{|\log\epsilon_{k}|}\inf_{{\rm diam}(\xi)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)>\overline{\rm mdim}_{M}(f,X,d)-\gamma.$
Let $\rho_{k}^{\ast}=\phi_{\ast}\rho_{k}\in M_{f\times f}(X\times X)$. Then it is easy to check that for every $k$,
@@
Indeed, let $\xi$ be a finite Borel partition of $X$, then $\eta=\{A\times B:A,B\in\xi\}$ is a partition of $X\times X$ satisfying ${\rm diam}(\eta,d\times d)={\rm diam}(\xi,d)$. For any $n\geq 1$,
@@
Let $K=\overline{{\rm Conv}\{\mu_{0},\mu_{1},\rho_{k}^{\ast},k\geq 1\}}$, i.e., the convex closure of set $\{\mu_{0},\mu_{1}\}\cup\{\rho_{k}^{\ast}:k\geq 1\}$. Then $K$ is a closed convex subset of $M_{f\times f}(X\times X)$. For any $(x_{0},y_{0})\in G_{K}$, there are sequences of integers $\{t_{n}^{0}\}$ and $\{t_{n}^{1}\}$ with $\mathcal{E}_{t_{n}^{0}}(x_{0},y_{0})\to\mu_{0}$, $\mathcal{E}_{t_{n}^{1}}(x_{0},y_{0})\to\mu_{1}$ as $n\to\infty$. Hence
@@
$=$ $\limsup_{\epsilon\to 0}\frac{\sup_{\nu\in K}\inf_{{\rm diam}(\eta)<\epsilon}h_{\nu}(f\times f,\eta)}{|\log\epsilon|}\geq\limsup_{k\to\infty}\frac{\inf_{{\rm diam}(\eta)<\epsilon_{k}}h_{\rho_{k}^{\ast}}(f\times f,\eta)}{|\log\epsilon_{k}|}$
$=$ $\limsup_{k\to\infty}\frac{\inf_{{\rm diam}(\xi)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)}{|\log\epsilon_{k}|}>\overline{\rm mdim}_{M}(f,X,d)-\gamma.$
Since $\gamma$ is arbitrary, one has
$\overline{\rm mdim}^{P}_{M}(f\times f,{\rm MLY}(f),d\times d)\geq\overline{\rm mdim}_{M}(f,X,d).$
The case of $\overline{\rm mdim}_{M}(f,d)=\infty$ is still valid by considering a slight modification. ∎
@@ 5 Examples
Applying the main results, we give the specific formula of the metric mean dimensions of generic points of full shift over compact metric space.
@@
Endow the product space $X=[0,1]^{\mathbb{Z}}$ with product metric
and let $\sigma:X\to X$ be the left shift map defined by $(x_{n})_{n\in\mathbb{Z}}\mapsto(x_{n+1})_{n\in\mathbb{Z}}$. It is shown that the system $(X,d,\sigma)$ has the specification property (cf. [DGS76, Proposition 21.2]), and ${\rm\overline{mdim}}_{M}(\sigma,X,d)=1$ [LT18]. Let $\mathcal{L}$ be the Lebesgue measure on $[0,1]$ and $\mu=\mathcal{L}^{\otimes\mathbb{Z}}$ is the product measure, which is $\sigma$-ergodic. Then the ergodicity enables us to define the upper $\epsilon$-Brin-Katok local entropy of $\mu$ as
$h_{\mu}^{BK}(\epsilon):=\limsup_{n\to\infty}-\frac{1}{n}\log\mu(B_{n}(x,\epsilon)),\ \mu\text{-a.e. }x\in X.$
@@
$\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi)=1={\rm\overline{mdim}}_{M}(\sigma,X,d).$
Then by Corollary 1.5, we have
@@
Let $K$ be a compact metric space with metric $d$. Consider the left shift map $\sigma$ on $K^{\mathbb{Z}}$ again, where $K^{\mathbb{Z}}$ is endowed with following product metric
@@
where $\overline{\rm dim}_{B}(F)$ is the upper box dimension of $F$ under the metric $d^{q}$.
@@
Let $K$ be a compact metric space with metric $d$. There exists $\mu\in M_{\sigma}(K^{\mathbb{Z}})$ such that
$\overline{\rm mdim}_{M}^{B}(\sigma,G_{\mu},d^{\mathbb{Z}})$ $=\overline{\rm mdim}_{M}^{P}(\sigma,G_{\mu},d^{\mathbb{Z}})=\overline{\rm mdim}_{M}^{UC}(\sigma,G_{\mu},d^{\mathbb{Z}})=\overline{\rm dim}_{B}(F)$
$=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}\limits h_{\mu}(\sigma,\xi).$
@@
$\hat{\mathcal{S}_{n}}=\{x\in K^{n}:x_{k}\in S_{\epsilon},\forall~{}0\leq k\leq n-1\},$
and consider the set $\mathcal{S}_{n}\subset X$ consisting of one preimage of each element of $\hat{\mathcal{S}_{n}}$ under $\pi_{n}$. Then $\hat{\mathcal{S}}_{n}$ is an $(n,\epsilon)$-separated set of $K^{\mathbb{Z}}$ and #$\#{\hat{\mathcal{S}}_{n}}=(\#S_{\epsilon})^{n}$. Define
@@
We assume that $\mu_{n}\rightarrow\mu_{\epsilon}\in M_{\sigma}(X)$. Then by Lemma 3.1, for each $\epsilon>0$ we have
@@
Assume that $\mu_{\epsilon}\rightarrow\mu\in M_{\sigma}(K^{\mathbb{Z}})$. Combining with Theorem 1.3, one has
@@
Finally, by Corollary 1.5 we complete the proof.
@@
Acknowledgments: The first and second authors were supported by NNSF of China (12101340). The third author was supported by the China Postdoctoral Science Foundation (Grant Number: 2024M763856); R.Yang would like to thank Yunping Wang for many helpful communication and the warm hospitality when he was visiting the Ningbo University.

### v2

@@ 4.1 Variational principle of level sets
denote the level set of $\alpha$ w.r.t. $\varphi$ consisting of the set of points of $X$ such that $\varphi$ is taking constant value $\alpha$ on $V_{f}(x)$. Then the continuity of $\varphi$ implies that $M_{f}(\varphi,\alpha)$ is a nonempty closed subset of $M_{f}(X)$.
Condition A: Given a continuous function $\varphi$ on $M_{f}(X)$, for every $\alpha\in L_{\varphi}$ we always assume that $M_{f}(\varphi,\alpha)$ is a convex subset.
@@
    Let $\Phi:=\{\phi_{n}\}$ be an asymptotically additive potential, i.e. for any $\delta>0$, there exists $\psi_{\delta}\in C(X)$ such that $\limsup_{n\to\infty}\frac{1}{n}||\phi_{n}-S_{n}\psi_{\delta}||<\delta$. We define
@@
    Specially, when the sequence $\{\phi_{n}\}$ is an additive function $\{S_{n}\psi\}$ with $\psi\in C(X)$, in this case we have
    $K_{\alpha}=\{x\in X:\forall\mu\in V_{f}(x),\ \int\psi\mathop{}\!\mathrm{d}\mu=\alpha\},$
    $M_{f}(\varphi,\alpha)=\{\mu\in M_{f}(X):\int\psi\mathop{}\!\mathrm{d}\mu=\alpha\}.$
Given a nonempty open subset $U\subset X$, we put
@@
Let $(X,d,f)$ be a dynamical system with specification property and $U\subset X$ be a nonempty open subset. Suppose that $\varphi:M_{f}(X)\to\mathbb{R}$ is a continuous function satisfying Condition A. Then for any $\alpha\in L$ and $S\in\{B,P\}$, we have
@@
$=$ $\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
@@
Put $G_{\mu}^{\omega,U}:=G_{\mu}\cap U\cap\left\{x\in X:\omega_{f}(x)=X\right\}$. By (3.14), there exists a constant $c>0$, which is not dependent on $\mu$, such that
@@
denote the $\alpha$-level set of $\psi$. If $\alpha\in\mathbb{R}$ is a constant such that $E_{\alpha}\not=\emptyset$, then
$\overline{\rm mdim}_{M}^{B}(f,E_{\alpha},d)=\overline{\rm mdim}_{M}^{P}(f,E_{\alpha},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{f}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
@@
$K_{\alpha}=\{x\in X:\forall\mu\in V_{f}(x),\int\psi\mathop{}\!\mathrm{d}\mu=\alpha\}.$
Fix $x\in X$. Since the statement $\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\psi(f^{j}x)=\alpha$ is equivalent to the statement $V_{f}(x)\subset M_{f}(\psi,\alpha)$, this shows $K_{\alpha}=E_{\alpha}$. We get the desired result by Theorem 4.1. ∎
Readers can turn to [O03, Tho09, ZC13] and references therein for more examples about the multifractal analysis of level sets using topological pressure. Then, by Theorem 4.1, one can get analogous variational principles for the metric mean dimensions of level sets of some special functions, which are determined by the dynamical behaviors of the system.
@@ 4.2 Linking metric mean dimension and mean Li-Yorke chaotic
In this subsection, we link the concepts of metric mean dimension and mean Li-Yorke chaos in the sense of infinite entropy. We give a quantitative characterization for the chaotic phenomenon of infinite entropy systems. Namely, we show the packing metric mean dimension of the set of mean Li-Yorke pairs is bounded from blow by the metric mean dimension of the phase space (Theorem 4.3).
@@
Let $(X,d,f)$ be a dynamical system with specification property. Then
$\overline{\rm mdim}^{UC}_{M}(f\times f,{\rm MLY}(f),d\times d)\geq\overline{\rm mdim}_{M}(f,X,d),$
$\overline{\rm mdim}_{M}^{UC}(f\times f,{\rm MLY}(f),d\times d)=\overline{\rm mdim}_{M}(f\times f,X\times X,d\times d).$
@@
Specification property implies that the system has positive entropy. Then by the well-known fact that positive entropy implies mean Li-Yorke chaotic (cf.[Dow14]), we can choose two distinct points of $x,y\in X$, a sequence $\{n_{i}\}$ of positive integer that converges to $\infty$ as $i\to\infty$, and $\theta>0$ (depending on $x,y$) such that $\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}d(f^{j}x,f^{j}y)>\theta$ for all sufficiently large $i$. We put $\mu_{n_{i}}=\frac{1}{n_{i}}\sum_{j=0}^{n_{i}-1}\delta_{(f^{j}x,f^{j}y)}$. Then any an accumulation point $\mu_{0}\in M_{f\times f}(X\times X)$ of $\{\mu_{n_{i}}\}$ satisfies the desired claim. ∎
Now take $\nu\in M_{f}(X)$ and put $\mu_{1}=\phi_{\ast}\nu\in M_{f\times f}(X\times X)$, where $\phi:X\to X\times X$ is the diagonal map given by $\phi(x)=(x,x)$. Then one has $\int_{X\times X}d(x,y)\mathop{}\!\mathrm{d}\mu_{1}=0$. Assume that $\overline{\rm mdim}_{M}(f,X,d)$ is finite. For any $\gamma>0$, by Theorem 2.4 we pick $\epsilon_{k}\to 0$ as $k$ goes to $\infty$ and $\rho_{k}\in M_{f}(X)$ such that for each $k$,
$\frac{1}{|\log\epsilon_{k}|}\inf_{{\rm diam}(\xi)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)>\overline{\rm mdim}_{M}(f,X,d)-\gamma.$ (4.4)
Let $\rho_{k}^{\ast}=\phi_{\ast}\rho_{k}\in M_{f\times f}(X\times X)$. Then for every $k$,
@@
Indeed, let $\xi$ be a finite Borel partition of $X$, then $\eta=\{A\times B:A,B\in\xi\}$ is a partition of $X\times X$ so that ${\rm diam}(\eta,d\times d)={\rm diam}(\xi,d)$. For any $n\geq 1$,
@@
Let $K=\overline{{\rm Conv}\{\mu_{0},\mu_{1},\rho_{k}^{\ast},k\geq 1\}}$, i.e. the convex closure of set $\{\mu_{0},\mu_{1}\}\cup\{\rho_{k}^{\ast}:k\geq 1\}$. Then $K$ is a closed convex subset of $M_{f\times f}(X\times X)$. For any $(x_{0},y_{0})\in G_{K}$, there are two sequences of integers $\{t_{n}^{0}\}$ and $\{t_{n}^{1}\}$ with $\mathcal{E}_{t_{n}^{0}}(x_{0},y_{0})\to\mu_{0}$, $\mathcal{E}_{t_{n}^{1}}(x_{0},y_{0})\to\mu_{1}$ as $n\to\infty$. Hence
@@
$=$ $\limsup_{\epsilon\to 0}\frac{\sup_{\nu\in K}\inf_{{\rm diam}(\eta)<\epsilon}h_{\nu}(f\times f,\eta)}{|\log\epsilon|}$
$\geq$ $\limsup_{k\to\infty}\frac{\inf_{{\rm diam}(\eta)<\epsilon_{k}}h_{\rho_{k}^{\ast}}(f\times f,\eta)}{|\log\epsilon_{k}|}$
$=$ $\limsup_{k\to\infty}\frac{\inf_{{\rm diam}(\xi)<\epsilon_{k}}h_{\rho_{k}}(f,\xi)}{|\log\epsilon_{k}|}\geq\overline{\rm mdim}_{M}(f,X,d)-\gamma.$
Since $\gamma$ is arbitrary, by Proposition 2.1 one has
$\overline{\rm mdim}^{UC}_{M}(f\times f,{\rm MLY}(f),d\times d)\geq\overline{\rm mdim}_{M}(f,X,d).$
The case of $\overline{\rm mdim}_{M}(f,X,d)=\infty$ is still valid by considering a slight modification.
For sufficiently small $\gamma>0$, let $c_{\gamma}:=\min\{\frac{1}{\gamma},\overline{\rm mdim}_{M}(f\times f,X\times X,d\times d)-\gamma\}$. By Theorem 2.4, we pick $\epsilon_{k}\to 0$ as $k$ goes to $\infty$ and $m_{k}\in M_{f\times f}(X\times X)$ such that for each $k$,
$\frac{1}{|\log\epsilon_{k}|}\inf_{{\rm diam}(\xi,d\times d)<\epsilon_{k}}h_{m_{k}}(f,\xi)>c_{\gamma}.$ (4.6)
Put $\hat{K}=\overline{{\rm Conv}\{\mu_{0},\mu_{1},m_{k},k\geq 1\}}$. Then the similar argument gives us
$\overline{\rm mdim}_{M}^{P}(f\times f,{\rm MLY}(f),d\times d)\geq\overline{\rm mdim}_{M}(f\times f,X\times X,d\times d),$
which implies the desired result by Proposition 2.1.
Finally, we give an example to verify the Corollary 4.2 and Theorem 4.3.
Example 4.4.
Endow the product space $[0,1]^{\mathbb{Z}}$ with the product metric
and let $\sigma:[0,1]^{\mathbb{Z}}\to[0,1]^{\mathbb{Z}}$ be the left shift map defined by $\sigma((x_{n})_{n\in\mathbb{Z}})\mapsto(x_{n+1})_{n\in\mathbb{Z}}$. It is shown that the system $([0,1]^{\mathbb{Z}},d,\sigma)$ has the specification property (cf. [DGS76, Proposition 21.2]), and ${\rm{mdim}}_{M}(\sigma,[0,1]^{\mathbb{Z}},d)=1$ [LT18].
  • (1)
    Thus, by the product formula of metric mean dimension one has ${\rm mdim}_{M}(\sigma\times\sigma,[0,1]^{\mathbb{Z}}\times[0,1]^{\mathbb{Z}},d\times d)=2$. Consequently, by Theorem 4.3, we have
    $\overline{\rm mdim}_{M}^{UC}(\sigma\times\sigma,{\rm MLY}(\sigma),d\times d)=2.$
  • (2)
    Let $\varphi:[0,1]^{\mathbb{Z}}\to\mathbb{R}$ be map defined by $\varphi((x_{n})_{n\in\mathbb{Z}})=x_{0}$. Then for any $\alpha\in[0,1]$, we have
    $\overline{\rm mdim}_{M}^{B}(\sigma,E_{\alpha},d)=\overline{\rm mdim}_{M}^{P}(\sigma,E_{\alpha},d)=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\sup_{\mu\in M_{\sigma}(\varphi,\alpha)}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi),$
[…]
@@ 5 Examples
Applying the main results, we give the specific formulae of the metric mean dimensions of generic points of full shift over compact metric space.
@@
Consider the Example 4.4 again. Let $\mathcal{L}$ be the Lebesgue measure on $[0,1]$ and $\mu=\mathcal{L}^{\otimes\mathbb{Z}}$ be the product measure, which is a $\sigma$-ergodic measure on $[0,1]^{\mathbb{Z}}$. Then the ergodicity of $\mu$ enables us to define the upper $\epsilon$-Brin-Katok local entropy of $\mu$ as
$h_{\mu}^{BK}(\epsilon):=\limsup_{n\to\infty}-\frac{1}{n}\log\mu(B_{n}(x,\epsilon)),\ \mu\text{-a.e. }x\in[0,1]^{\mathbb{Z}}.$
@@
$\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi)=1={\rm\overline{mdim}}_{M}(\sigma,[0,1]^{\mathbb{Z}},d).$
Then, by Corollary 1.5, we have
@@
Let $K$ be a compact metric space with a metric $d$. Consider the left shift map $\sigma$ on $K^{\mathbb{Z}}$ again, where $K^{\mathbb{Z}}$ is endowed with following product metric
@@
where $\overline{\rm dim}_{B}(K)$ is the upper box dimension of $K$ under the metric $d$.
@@
Let $K$ be a compact metric space with a metric $d$. Then there exists $\mu\in M_{\sigma}(K^{\mathbb{Z}})$ such that
$\overline{\rm mdim}_{M}^{B}(\sigma,G_{\mu},d^{\mathbb{Z}})$ $=\overline{\rm mdim}_{M}^{P}(\sigma,G_{\mu},d^{\mathbb{Z}})=\overline{\rm mdim}_{M}^{UC}(\sigma,G_{\mu},d^{\mathbb{Z}})=\overline{\rm dim}_{B}(K)$
$=\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(\sigma,\xi).$
@@
${\mathcal{S}_{n}}=\{x\in K^{n}:x_{k}\in S_{\epsilon},\forall~{}0\leq k\leq n-1\},$
and consider the set $\hat{\mathcal{S}}_{n}\subset K^{\mathbb{Z}}$ consisting of one preimage of each element in ${\mathcal{S}_{n}}$ under $\pi_{n}$. Then $\hat{\mathcal{S}}_{n}$ is an $(n,\epsilon)$-separated set of $K^{\mathbb{Z}}$ and #$\#{\hat{\mathcal{S}}_{n}}=(\#S_{\epsilon})^{n}$. Define
@@
We assume that $\mu_{n}\rightarrow\mu_{\epsilon}\in M_{\sigma}(K^{\mathbb{Z}})$. Then, by Lemma 3.1, for each $\epsilon>0$ we have
@@
Assume that $\mu_{\epsilon}\rightarrow\mu\in M_{\sigma}(K^{\mathbb{Z}})$ as $\epsilon\to 0$. Combining with the Theorem 1.3, one has
@@
Finally, by the Corollary 1.5 we complete the proof.
@@
Acknowledgments: The first and second authors were supported by NNSF of China (12101340). The third author was supported by the China Postdoctoral Science Foundation (Grant Number: 2024M763856); R.Yang would like to thank Yunping Wang for many helpful communication and the warm hospitality when he was visiting the Ningbo University. The authors also thank Prof. Ercai Chen for many useful comments.